OpenStudy (anonymous):

Consider the following polygons: file:///C:/Documents%20and%20Settings/name/My%20Documents/IMG_20140615_234705.jpg file:///C:/Documents%20and%20Settings/name/My%20Documents/IMG_20140615_234711.jpg file:///C:/Documents%20and%20Settings/name/My%20Documents/IMG_20140615_234717.jpg

OpenStudy (anonymous):

Find the area of each of the polygons. Suppose each side of each of the polygons is doubled. Find the new area. Suppose each side of each of the polygons is tripled. Find the new area. What conclusions can you make about how the polygon’s area as their size increases proportionately? How would the area increase if each side of each polygon is quadrupled?

OpenStudy (neer2890):

upload your files pls?

OpenStudy (neer2890):

or images?

OpenStudy (neer2890):

\[Area of \triangle=\frac{ Base*Altitude }{ 2 }\]

OpenStudy (neer2890):

so now solve for 1st question and tell me?

OpenStudy (neer2890):

Base=4 m and Altitude=4m. put them in formula and find area?

OpenStudy (anonymous):

I don't really know to do that.

OpenStudy (neer2890):

i've given you the formula In 1st figure, Base of triangle= 4m and altitude or height= 4m \[Area=\frac{ 4*4 }{ 2 }=\frac{ 16 }{ 2 }=8 m ^{2}\]

OpenStudy (neer2890):

ok, now if the sides of triangle are doubled. i.e., Base= 8m, Height=8m and Hypotenuse=12m now tell me the area of new triangle formed?

OpenStudy (neer2890):

but there is a problem with your question, your triangle is not right angled as \[4^{2}+4^{2} \neq 6^{2}\]

OpenStudy (anonymous):

oh yes on the first pic it's a 6

OpenStudy (neer2890):

so (4,4,6) is not a pythagoras triplet.

OpenStudy (neer2890):

Coming to 2nd figure. Area of Square= side*side(or square of side as all sides of square are equal). so side of your square=4m. so Area= 4*4= 16 m^2.

OpenStudy (neer2890):

now double the sides of your square. i.e., side = 8m( all sides are of 8m length.) now find the area and tell me?

OpenStudy (neer2890):

for 3rd figure, Area of rectangle= Length*Breadth Length=10 m and Breadth= 4m so Area= 10*4= 40 m^2

OpenStudy (neer2890):

now double the sides of your rectangle and find the area?